# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v4

@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. The current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting.
To address this issue, we propose our continual dynamic neural graphics primitives (CD-NGP) based on parameter isolation. We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly. The $k$-th branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects. Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals. We thus substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting as:
@@
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as Eq. 4 with the color MLP $f_{\theta_{2k}}$. For simplicity, we denote the models adopting pure voxel encoders in Eq. 5 as CD-NGP, the models adopting the hybrid encoders in Eq. 6 as CD-MERF.
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most points ($>90\%$) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}$. Please note that $P_{1}$ is significantly larger than $P_{2}$ to ensure the scalability of the model. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our assymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
2.4 Rendering and optimization
$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (8)
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray. The model is optimized by L2 photometric loss with three regularization terms, defined as:
$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r}$ (9)
$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})$ (10)
$L_{o}$ in Eq. 9 is the entropy on the opacity by [25], defined as:
$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j}.$ (11)
$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (12)
We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
2.5 Continual learning pipeline.
We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches. The whole model is trained chunk by chunk. The implementation will be available upon publication.
@@ 2.4 Rendering and optimization
0:  $N_{chunk},T_{chunk},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
@@ 3.1 Datasets
Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution. The lengths of the scenes are set as 300 frames and the refresh rates are set as 30 FPS.
To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$. Similar to the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. We synchronize the multi-view videos using environmental sound and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation.
The dynamics in our dataset are notably more challenging than those in the DyNeRF dataset, featuring significant non-rigid motions of human bodies and transient effects such as slide effects on the screens in the backgrounds. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset.
This dataset was meticulously collected to evaluate the scalability of our method and to support further research. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants. The view in the upper center is designated for validation.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq. 6 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 3.
At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
3.3 Results and comparisons
Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset. Memory denotes the maximum host memory required for training. - denotes the metrics unreported in the original publications. Best, Second, and Third metrics are in highlight. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@ 3.3 Results and comparisons
CD-NGP (Ours) $\checkmark$ 10 30.23 0.029 0.198 75 min 113MB 14GB
3.3.1 DyNeRF dataset.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset. StreamRF [27] and INV [48] are trained per-frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
@@
We compare our method with both online and offline baselines. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
@@
Figure 3 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based continual learning methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4.
@@
3.3.2 Meetroom dataset.
Table 2 and Fig. 5 shows the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6 is the PSNR curve of the compared methods on the trimming scene. As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig. 6. Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig. 5. We use the released configuration of StreamRF [27] for the meetroom dataset. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
Figure 6: PSNR curves of the compared methods on the trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively.
Table 2: Results on the Meetroom dataset.
@@
Mixvoxels $\checkmark$ 26.96 0.027 0.1034 15min 500MB
HexPlane $\checkmark$ 24.71 0.043 0.2206 100min 200MB
HexPlane-GR $\checkmark$ 19.05 0.099 0.3999 100min 318MB
HexPlane-PR $\checkmark$ 18.63 0.112 0.3589 100min 318MB
StreamRF $\checkmark$ 26.48 0.029 0.1699 75min 1710MB
INV $\checkmark$ 23.87 0.041 0.1195 40hours 336MB
CD-NGP (Ours) $\checkmark$ 26.01 0.039 0.1906 77min 113MB
3.3.3 Spatial representations.
Table 3: Ablation study on spatial representations. Different representations lead to subtle changes in quality metrics.

### v5

@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
As defined in Section 2.1, we split the model into $N_{chunk}$ individual branches to represent the $N_{chunk}$ chunks of the dynamic scene. Consequently, the spatial features encoded by different model branches could be fused to reduce the number of the parameters and improve scalability. This approach, defined as continual dynamic neural graphics primitives (CD-NGP), is designed to handle the dynamic nature of scenes in the online setting. Equipped with the hash encoders and MLP regressors defined in Section 2.2, the $k$-th branch of the model contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. Please note the dimensions of the features is independent of the sizes of the hash tables: a larger hash table alleviates hash collision at the cost of more storage consumption, but the features encoded by different sizes of hash tables share the same size, i.e. $\Psi_{k}(x)\in\mathbf{R}^{LF}$ holds for all $k$. To alleviate hash collision and reduce model size simultaneously, we apply an asymmetric hash table design: the hash table in the first branch (base branch) is set as $2^{P_{1}}$, and the hash tables in the subsequent branches (auxiliary branches) are set as $2^{P_{2}}$, where $P_{1}$ is much larger than $P_{2}$. The base branch is designed to reconstruct the scenes with initial movements and static objects, and the auxiliary branches are designed to compensate for residuals. The re-use of the features in the base branch ensures high scalability and alleviates the catastrophic forgetting problem. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. We substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting, defined as:
@@
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as in Eq. 4 with the color MLP $f_{\theta_{2k}}$. To cope with large motions and complex transient effects in the long time intervals, we assign $T_{episode}$ chunks as an episode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. We further illustrate the whole process of our method in Algorithm 1, where $\eta_{init}$ and $\eta_{aux}$ denote the numbers of iterations of the initial branch and the auxiliary branches, respectively. $\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)}$ denote the model parameters in the branches.
@@
0:  $N_{chunk},T_{chunk},T_{episode},\eta_{init},\eta_{aux},\Psi_{i},\Gamma_{i},\theta_{1}^{(i)},\theta_{2}^{(i)},i\in\mathbb{N}$
@@
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most points ($>90\%$ in DyNeRF dataset by their observation) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our asymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.
2.4 Rendering and optimization
$\hat{C}(r)$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c_{i}}=\sum_{i=1}^{N}T_{i}(1-e^{-\sigma_{i}\delta_{i}})\mathbf{c_{i}},$ (7)
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ denotes the transmittance, $\delta_{i}$ is the interval between neighboring samples, $N$ is the number of samples along the ray, $\mathbf{c_{i}}$ is the output of Eq. 4. The model is optimized by L2 photometric loss with three regularization terms, defined as:
$L=\frac{1}{|\mathbf{R}|}\sum_{r\in\mathbf{R}}\|C(r)-\hat{C}(r)\|_{2}^{2}+\lambda_{d}L_{d}+\lambda_{o}L_{o}+\lambda_{r}L_{r},$ (8)
$+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i}),$ (9)
$L_{o}$ in Eq. 8 is the entropy on the opacity [25], defined as:
$L_{o}=-o\log(o),\quad o=\sum_{j=1}^{N}T_{j}\alpha_{j},$ (10)
$L_{r}=||\Psi_{k}(\mathbf{x})||_{1}.$ (11)
We set $\lambda_{r}$ as 0.001 to encourage using the features encoded by the base branch. Based on [25], we set a large value of $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3.1 Datasets
Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution. The scenes are 300 frames long and captured at 30 FPS.
To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig. 2. Similar to those in the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. The cameras are arranged in two rows and placed in an arc facing the scenes. The camera in the center of the upper row is designated for validation. We synchronize the multi-view videos and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation. Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset. Our dataset is collected to evaluate the scalability of our method and to support further research in the community. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following most of the state-of-the-art methods [28, 27, 17, 45, 43], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We follow [27] and use $1280\times 720$ resolution on the Meetroom dataset. We use $1280\times 720$ resolution for our dataset to fit into the host memory. We use the importance sampling approach described in [28] at a batch size of 1024 rays. The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability. The learning rate is set as 0.001 and scheduled by cosine function. We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches. The structural parameters $(L,F)$ in the spatial hash encoders are set as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for $\eta_{init}=18k$ iterations, and for each of the subsequent branches we train for $\eta_{aux}=3k$ iterations using the previous branch for initialization for the DyNeRF dataset and our dataset. We use $\eta_{init}=33k$ and $\eta_{aux}=3k$ for the Meetroom dataset. We use LeakyReLU activation in the MLP regressors. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our own dataset. We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene.
3.3 Results and scalability analysis
Table 1: Comparisons between our CD-NGP and state-of-the-art methods on the DyNeRF dataset. Memory denotes the maximum host memory required during training. $\dagger$ denotes the methods evaluated only on the flame salmon scene. - denotes the metrics not reported in the original publications.
@@ 3.3 Results and scalability analysis
CD-NGP (ours) $\checkmark$ 10 30.23 0.027 0.198 75 min 113MB 14GB
3.3.1 Comparisons on DyNeRF dataset.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. The methods in [28, 34] are trained on 8 V100 GPUs or equivalently 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. We report the metrics in the publication if not otherwise specified. HexPlane [7] uses resolution $1024\times 768$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset. StreamRF [27] and INV [48] are trained per frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR for pixel-wise similarities, DSSIM for structural similarities, and LPIPS [57] for perceptual similarities (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
We compare our method with both online and offline baselines. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also reduces $40\%$ less model size (to $0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches and delivers comparable quality to the offline methods.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based online methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4.
@@
3.3.2 Comparisons on the Meetroom dataset.
Table 2 and Fig. 5 show the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6(a) is the PSNR curve of the compared methods on the trimming scene. We use the released configuration of StreamRF [27] for the meetroom dataset to obtain the results. As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the meetroom dataset, we thus exclude them from Fig. 6(a) for better visualization. The metrics in Table 2 and the curves in Fig. 6(a) show the advantage of our method over both online and offline baselines. It outperforms the implicit MLP-based online method [48] regarding reconstruction quality, speed, and model size. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
Figure 6: Combined PSNR curves for various scenes and methods. (a) The trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively. (b) The first 1200 frames of seminar scene in our dataset. The solid lines and dotted lines denote offline and online methods respectively. (c) The first 2400 frames of seminar scene in our long video dataset. (d) the sear steak scene in the DyNeRF dataset. Solid and dotted lines denote symmetric and asymmetric setting, respectively.
Table 2: Comparisons of the metrics on the Meetroom dataset.
@@
Mixvoxels $\checkmark$ 26.96 0.027 0.103 15min 500MB
HexPlane $\checkmark$ 24.71 0.043 0.221 100min 200MB
HexPlane-PR $\checkmark$ 18.63 0.112 0.359 100min 318MB
HexPlane-GR $\checkmark$ 19.49 0.099 0.400 100min 318MB
StreamRF $\checkmark$ 26.48 0.029 0.170 75min 1710MB
INV $\checkmark$ 23.87 0.041 0.120 40hours 336MB
CD-NGP (ours) $\checkmark$ 26.01 0.039 0.191 70min 113MB
3.3.3 Spatial representations and scalability analysis.
Our proposed continual learning framework is robust against changes in spatial backbones, including spatial hash encoder configurations and feature fusion methods. The hash encoder in Eq. 5 could be changed into the hybrid representation (MERF) [38] and the plane representation like those in [17, 21]. The MERF representation is defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x),$ (12)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated features from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m$-th resolution. The plane representation is defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x).$ (13)
For the simplicity of the notations, we denote the models adopting Eq. 5, Eq. 12, and Eq. 13 as CD-NGP, CD-MERF, and CD-Plane respectively.
We compare the results of different spatial representations on the DyNeRF dataset and list the parameter numbers in Table 3. The alternatives of spatial representations include CD-MERF and CD-Plane. The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments, we thus remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. As shown in Table 3, the hybrid representation CD-MERF achieves high compression and comparable model sizes to methods for static scenarios [31, 21] ($\sim$60MB) without significant quality loss. We also adopt the 3D+4D representation similar to that in [34] for comparison. This representation is implemented by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5, i.e. $\Gamma(x,t)=\bigoplus_{m=1}^{L}P_{m}(x,y,z,t)$. However, the asymmetric hash sizes (using $(P_{1},P_{2})=(19,14)$) in the 3D+4D representation results in serious feature collisions and deliver bad results (PSNR < 16 dB on 3 scenes). We thus set $(P_{1},P_{2})=(14,14)$ for the 3D+4D representation in Table 3 to keep the size of the model close to our CD-NGP for fair comparison. The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
The results in Table 3 also show the flexibility in the feature fusion methods. We replace the element-wise sum in Eqs. 6 and 12 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3. Benefiting from larger feature dimensions, these methods lead to slightly better quality (especially when measured in LPIPS) at the cost of more training time, and the reasonable trade-off also shows the robustness of our method.
Table 3: Quality metrics, parameter numbers, and bandwidth costs of different spatial representations. The parameter numbers of each module of the 30-branch representations are also shown. #$\#3DEnc.$ and #$\#2DEnc.$ denote the parameter numbers of the 3D encoders and 2D encoders in the base branch and one auxiliary branch accordingly. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).